# Gravitational waves from cosmic strings in LISA: reconstruction pipeline and physics interpretation

### v1

@@ 1 Introduction
In this paper, we will focus on models in which loops decay by emitting GWs only, corresponding to the maximal SGWB. This corresponds to assuming that cosmic strings may be well described by the Nambu-Goto action or, in order words, that cosmic strings have no internal DoF and their thickness $\delta\sim 1/\eta$ is much smaller than their typical curvature length. Field theory cosmic strings, however, may also decay into excitations of their constituent fields, which could result in a weakening of their GW signals. For Abelian-Higgs strings — local strings that are generally expected to be well described by the Nambu-Goto action — the potential role of these other decay mechanisms is still a matter of debate Vachaspati:1984yi ; Vincent:1997cx ; Moore:1998gp ; Olum:1998ag ; Olum:1999sg ; Moore:2001px ; Hindmarsh:2008dw ; Hindmarsh:2017qff ; Matsunami:2019fss ; Hindmarsh:2021mnl ; Blanco-Pillado:2022rad ; Hindmarsh:2022awe ; Blanco-Pillado:2023sap . Although in numerical simulations of the whole network loops decay quickly into scalar and gauge radiation Hindmarsh:2017qff , individual loop simulations indicate that this radiation is mostly emitted in high-curvature regions Olum:1998ag ; Olum:1999sg ; Matsunami:2019fss or in the end stage of their evolution, when loop length becomes comparable to $\delta$. This phenomenon has also been recently observed in individual loops extracted directly from cosmological field theory simulations of networks of strings Blanco-Pillado:2023sap . This gives more credence to the idea that indeed field theory networks lose their energy from these high curvature regions. Here, since we consider strings $\delta\lesssim 10^{-24.5}\,\text{cm}$ and loops of astrophysical sizes, we will generally neglect the energy lost as a result of these other decay mechanisms. We only relax this assumption in appendix C where we consider situations in which only a part of the energy of loops goes into GWs and/or in which only a fraction of the loops decay by emitting significant gravitational radiation.
@@
Our analysis, however, does not apply to other more complex cosmic string models, such as cosmic superstrings Sarangi:2002yt ; Dvali:2003zj ; Copeland:2003bj , metastable cosmic strings Preskill:1992ck ; Leblond:2009fq ; Buchmuller:2020lbh ; Buchmuller:2023aus ; Chitose:2023dam , superconducting strings Witten:1984eb ; Nielsen:1987fy , or whose energy is not localized, such as global strings Vilenkin:1982ks . Superconducting string loops, for instance, carry current and are then expected to emit significant electromagnetic radiation Vilenkin:1986zz ; Blanco-Pillado:2000nbp and this current was shown to lead to a suppression of the gravitational radiation power emitted too Babichev:2003au ; Rybak:2022sbo . For global or axion strings, although an SGWB is inevitable and potentially detectable Chang:2019mza ; Chang:2021afa ; Gorghetto:2021fsn ; Figueroa:2020lvo , the emission of GWs by loops is, in general, subdominant to the emission of Goldstone bosons/axions which would result in significant suppression of the amplitude of this background Vilenkin:1986ku ; Drew:2019mzc ; Drew:2022iqz ; Baeza-Ballesteros:2023say ; Drew:2023ptp .
There are also alternative models of the loop number density that we will not consider here. For instance, refs. Ringeval:2005kr ; Lorenz:2010sm ; Auclair:2019zoz ; Auclair:2020oww propose a divergent loop production function with a cut-off at the gravitational backreaction scale. These loop number densities lead to an additional (very large) population of small loops that has a significant impact on the predicted GW spectrum Ringeval:2017eww ; LIGOScientific:2017ikf . However, the validity of such models has been recently put into question Blanco-Pillado:2019vcs ; Blanco-Pillado:2019tbi and we will not discuss them here  [^7] For an analysis including this alternative model see for example Boileau:2021gbr ..
@@
This paper is laid out as follows. In section 2, we review SGWB production by cosmic strings and discuss the two models used in this study. The reader familiar with cosmic string SGWB and the models used (Models I and II of ref. Auclair:2019wcv ) can likely skip much of this section, but may wish to stop at section 2.2, where we introduce a novel way to include all relevant emission modes in the computation of the template for Model I, and in section 2.4, where we discuss the effects of non-standard cosmological histories on the SGWB. In section 3, we go into further details about how we implement the cosmic-string SGWB search and reconstruction in the SGWBinner. This involves reconstructing signals using templates and incorporating elements such as the LISA noise curve and GWs from other sources. In section 4, we use SGWBinner to forecast LISA’s ability to reconstruct the cosmic string SGWB for both models under various circumstances, most notably in the presence of astrophysical foregrounds from binary systems. Finally, in section 5, we comment on the physics interpretation of the potential results obtained by the LISA observatory. We conclude in section 6.
@@ 2.1 Review of SGWB production by cosmic strings
addresses the deviation from the naive $T\propto a^{-1}$ relation dictated by entropy conservation, and represents the dependence on the effective number of relativistic degrees of freedom (DoF): $g_{*}$ for energy and $g_{*S}$ for entropy. Here we will adopt the parametrization of the time evolution of $g_{*}$ and $g_{*S}$ in the Standard Model (SM) of particle physics given by micrOMEGAs_3.6.9.2 Belanger:2014vza . While it is possible to make other choices for the various inputs to the expansion history while remaining within “standard” cosmology, these choices have only a minor impact on the SGWB generated by cosmic string; see, however, sections 2.4, 4.3 for the effects of more significant changes to this standard history.
@@ 2.2 Template based on analytical modeling (Model I)
  • Loops are created with a length that is a fixed fraction $\alpha_{L}\leq 1$ of the characteristic length $L$ of the cosmic string network.
Under these assumptions, a description of the evolution of the energy density of the network is sufficient to fully characterize the loop distribution function throughout cosmic history and to compute the SGWB they generate. Here, as in refs. Sousa:2013aaa ; Sousa:2014gka , we use the Velocity-dependent One-scale Model Martins:1996jp to describe the evolution of the network.
In this model, the loop-size parameter $\alpha=\alpha_{L}L/t$, which characterizes the loop length at formation, is treated as a free parameter and a power spectrum of the form in eq. (3) is assumed. This model then has, besides cosmic string tension $G\mu$, two additional parameters: $\alpha$ and the spectral index $q$. Note that although the assumptions of the model may seem rather strong, they may be easily relaxed by assuming that only a fraction $\mathcal{F}$ of the energy lost by the network goes into the production of loops that decay by emitting GWs (as suggested, for instance, in ref. Hindmarsh:2022awe ). Such a parameter may also be used to describe, to some extent, situations in which loops are created with a distribution of lengths (and kinetic energies) instead of being created with exactly the same size Blanco-Pillado:2013qja ; Sousa:2020sxs   [^9] Note that the results obtained under these assumptions may be used to construct the SGWB for any distribution of loop length at the moment of creation, as described in refs. Sanidas:2012ee ; Sousa:2020sxs .. This fuzziness parameter, accounting for potential uncertainties in the modeling of cosmic string loops, may also be treated as a fourth free parameter of the model. Here, unless stated otherwise, we will take $\mathcal{F}=0.1$ as suggested by the results of Nambu-Goto simulations Blanco-Pillado:2013qja .
The contribution of the $j$-th harmonic mode of emission to the SGWB is such that $j\mathrm{\Omega}_{\rm gw}^{j}(jf)=\mathrm{\Omega}_{\rm gw}^{1}(f)$ (cf. eq. (4)) and therefore determining the contribution of the fundamental mode of emission is sufficient to fully characterize the SGWB generated by cosmic string loops. In ref. Sousa:2020sxs , the authors derived accurate analytical approximations to the three contributions $\Omega^{r}_{\rm gw}$, $\Omega^{rm}_{\rm gw}$, and $\Omega^{m}_{\rm gw}$ in the fundamental mode of emission. Here, these will serve as the basis to construct a fully analytical template for Model I, $\Omega_{\rm gw}(f,q,G\mu,\alpha)$, including all relevant modes of emission in summation in eq. (2). Since this derivation is quite lengthy and technical, here we just provide a brief outline of the method used. We refer the reader to ref. Sousa:2020sxs and appendix A for a detailed derivation.
@@
where $i=r$ for $\mathrm{\Omega}_{\rm gw}^{r}$ and $\mathrm{\Omega}_{\rm gw}^{rm}$ and $i=m$ for $\mathrm{\Omega}_{\rm gw}^{m}$. Nevertheless, since computing this summation explicitly is computationally costly, we resort to the Euler–Maclaurin summation formula knopp1990theory to perform it.
Here, we have also included the impact of the decrease of the number of relativistic DoF that occurs during the radiation era in $\mathrm{\Omega}_{\rm gw}^{r}$ (which was not originally considered). This decrease, encoded in the function $\Delta_{r}(a)$ in (7), affects the expansion rate of the universe and thus may leave clear signatures on the spectrum. Although the change in $\Delta_{r}(a)$ is continuous, here we model it as a piecewise constant function based on micrOMEGAS data Belanger:2014vza , and follow the approach in ref. Sousa:2020sxs to derive an analytical correction to $\mathrm{\Omega}_{\rm gw}^{r}$ to account for this effect. Note however that, when there is a sudden change in the rate of expansion caused by a change in the number of relativistic DoF, the evolution of the string network temporarily deviates from the linear scaling regime and the network takes some time to relax back into this regime. This effect is non-linear and complex to model analytically and, for this reason, we develop a fit using the full numerical computation of the SGWB (obtained using the approach in ref. Sousa:2013aaa ). We also describe this fit in more detail in appendix A.
@@ 2.3 Template based on simulation-inferred modeling (Model II)
We characterize $P_{j}$ by what is known as the BOS spectrum Blanco-Pillado:2015ana ; Blanco-Pillado:2017oxo , instead of using the power-law approximation in (3101010The literature often uses “Model II” to mean a cosmic string SGWB model which adopts the loop number densities of eq. (9) without specifying the $P_{j}$, i.e., without distinguishing between the power-law and BOS approaches. Here, by “Model II”, we will always mean the one which uses the BOS $P_{j}$. We will continue to specify this approach in the text to avoid any confusion with the literature. This spectrum is computed from a set of around $1000$ non-self-intersecting loops extracted from the scaling population of loops obtained from the cosmological network simulations Blanco-Pillado:2017oxo . The calculation of this spectrum takes into account backreaction by using a toy model based on a smoothing procedure. The result of this approach leads to a cusp-like spectrum at high $j$ and approximately matches with Model I when that model sets $\alpha=0.1$, $q=4/3$, and $\mathcal{F}=0.1$ (see Appendix A). The spectra are thus similar to those in the top left panel of figure 1  [^11] For a more comprehensive comparison of the spectra predicted in model I and model II we refer the reader to Auclair:2019wcv ..
Creating an SGWB using Model II involves the use of data tables, most notably when taking the BOS power spectrum, but also in the DoF $\Delta_{r}(a)$, which we take from micrOMEGAs Belanger:2014vza . As such, there is no closed-form expression for this $\Omega_{\rm gw}$ (even in the instance of analytic, single-index spectra such as the cusp or kink models). Instead, for the template subroutine, we make use of a comprehensive data table, plus interpolation, to determine the SGWB injected and recovered.
@@ 2.4 Effects of non-standard cosmologies on the SGWB
I. Modified equation of state. The first type of cosmological modification we consider is an early period, preceding the recent radiation dominated era, during which the universe is dominated by a new energy component, leading to a non-standard equation of state for the cosmos Allahverdi:2020bys . For example, an early matter domination era with, $\rho\propto a^{-3}$, may result from an appreciable abundance of a long-lived massive particle, oscillations of scalar fields or primordial black holes Moroi:1999zb ; Gouttenoire:2019rtn ; Bernal:2021yyb ; Ghoshal:2023sfa . Such an era ends with a reheating-like process as the long-lived species decay into SM particles. Another example is kination domination, where the cosmic energy is dominated by a scalar field evolving in a potential $V(\phi)\propto\phi^{N_{\phi}}$, which results in $\rho\propto a^{-n}$, with $n=6N_{\phi}/(N_{\phi}+2)$. In the limit of large $N_{\phi}$, $N_{\phi}\to\infty$, we have $n\to 6$ and the oscillation energy is dominated by the kinetic energy of the scalar. Kination may arise in certain theories related to inflation, quintessence, dark energy, and axions Salati:2002md ; Chung:2007vz ; Poulin:2018dzj ; Gouttenoire:2021jhk . While we will not consider these cases directly, more complicated histories including combinations of these mentioned above are also possible Gouttenoire:2021jhk ; Co:2021lkc ; Servant:2023tua . In order to retain the successful predictions of BBN theory Hannestad:2004px , these new phases in cosmic evolution must transit to the standard radiation domination before $T\sim 5~{}\rm MeV$.
Let us then assume that, before the recent radiation era, at temperatures $T>T_{\rm rd}$, the energy density of the universe is dominated by a component whose energy density scales as $\rho\propto a^{-3(1+w)}$, where $w$ is the equation-of-state parameter for this period. In this case, for large enough frequencies $f>f_{\rm rd}$, the contribution of the fundamental mode of emission to the SGWB is modified as  Cui:2017ufi ; Cui:2018rwi ; Sousa:2020sxs :

### v2

@@ 1 Introduction
In this paper, we will focus on models in which loops decay by emitting GWs only, corresponding to the maximal SGWB. This corresponds to assuming that cosmic strings may be well described by the Nambu-Goto action or, in order words, that cosmic strings have no internal degrees of freedom and their thickness $\delta\sim 1/\eta$ is much smaller than their typical curvature length. Field theory cosmic strings, however, may also decay into excitations of their constituent fields, which could result in a weakening of their GW signals. For Abelian-Higgs strings — local strings that are generally expected to be well described by the Nambu-Goto action — the potential role of these other decay mechanisms is still a matter of debate Vachaspati:1984yi ; Vincent:1997cx ; Moore:1998gp ; Olum:1998ag ; Olum:1999sg ; Moore:2001px ; Hindmarsh:2008dw ; Hindmarsh:2017qff ; Matsunami:2019fss ; Hindmarsh:2021mnl ; Blanco-Pillado:2022rad ; Hindmarsh:2022awe ; Blanco-Pillado:2023sap . For the majority of this work, we will consider only energy loss into gravitational waves, consistent with prior work within the LISA Consortium Auclair:2019wcv . We only relax this assumption in Sec. 4.5 where we consider situations in which only a part of the energy of loops goes into GWs, with the remainder being lost into other particle radiation.
@@
Our analysis, however, does not apply to other more complex cosmic string models, such as cosmic superstrings Sarangi:2002yt ; Dvali:2003zj ; Copeland:2003bj , metastable cosmic strings Preskill:1992ck ; Leblond:2009fq ; Buchmuller:2020lbh ; Buchmuller:2023aus ; Chitose:2023dam , superconducting strings Witten:1984eb ; Nielsen:1987fy , or models whose energy is not localized, such as global strings Vilenkin:1982ks . Superconducting string loops, for instance, carry current and are then expected to emit significant electromagnetic radiation Vilenkin:1986zz ; Blanco-Pillado:2000nbp and this current was shown to lead to a suppression of the gravitational radiation power emitted too Babichev:2003au ; Rybak:2022sbo . For global or axion strings, although an SGWB is inevitable and potentially detectable Chang:2019mza ; Chang:2021afa ; Gorghetto:2021fsn ; Figueroa:2020lvo , the emission of GWs by loops is, in general, subdominant to the emission of Goldstone bosons/axions which would result in significant suppression of the amplitude of this background Vilenkin:1986ku ; Drew:2019mzc ; Drew:2022iqz ; Baeza-Ballesteros:2023say ; Drew:2023ptp .
@@
Transient gravitational waves are also of significant interest theoretically and experimentally. LISA will be sensitive to many individual sources, including possible “bursts” from cosmic strings LISAConsortiumWaveformWorkingGroup:2023arg . We will not consider such bursts in this paper, as our focus is on SGWB.
@@
This paper is laid out as follows. In section 2, we review SGWB production by cosmic strings and discuss the two models used in this study. The reader familiar with cosmic string SGWB and the models used (Models I and II of ref. Auclair:2019wcv ) can likely skip much of this section, but may wish to stop at section 2.2, where we introduce a novel way to include all relevant emission modes in the computation of the template for Model I, and in section 2.4, where we discuss the effects of non-standard cosmological histories on the SGWB. In section 3, we go into further details about how we implement the cosmic-string SGWB search and reconstruction in the SGWBinner. This involves reconstructing signals using templates and incorporating elements such as the LISA noise curve and GWs from other sources. In section 4, we use SGWBinner to forecast LISA’s ability to reconstruct the cosmic string SGWB for both models under various circumstances, most notably in the presence of astrophysical foregrounds from binary systems. In section 5, we comment on the physics interpretation of the potential results obtained by the LISA observatory. We conclude in section 6. We finally devote appendix A and appendix B to some original, technical results on the templates investigated in our analysis.
@@ 2.1 Review of SGWB production by cosmic strings
addresses the deviation from the naive $T\propto a^{-1}$ relation dictated by entropy conservation, and represents the dependence on the effective number of relativistic degrees of freedom (DoF): $g_{*}$ for energy and $g_{*S}$ for entropy. Here we will adopt the parametrization of the time evolution of $g_{*}$ and $g_{*S}$ in the Standard Model (SM) of particle physics using the Gondolo–Gelmini equation of state GONDOLO1991145 as implemented in micrOMEGAs5.0 Belanger:2018ccd . While it is possible to make other choices for the various inputs to the expansion history while remaining within “standard” cosmology, these choices have only a minor impact on the SGWB generated by cosmic string; see, however, sections 2.4, 4.3 for the effects of more significant changes to this standard history.
The following two subsections outline two commonly-used approaches to constructing a string GWB template: one based on analytical modeling, and one based on simulation-inferred modeling. These models agree closely in their predictions for the shape and amplitude of resulting string GWBs, and thus in their predictions on LISA’s ability to detect or constrain strings. There is also a simulation-inferred model of the loop number density which has been included in LISA Consortium work (see Refs. Auclair:2019wcv ; LISACosmologyWorkingGroup:2022jok ), but which we do not consider here; we direct the reader to Refs. Ringeval:2005kr ; Lorenz:2010sm ; Auclair:2019zoz ; Auclair:2020oww ; Blanco-Pillado:2019vcs ; Blanco-Pillado:2019tbi ; Auclair:2021jud for details on this model. That model predicts a very large population of small loops that can lead to significantly different predictions for the GWB than in the models we consider here Ringeval:2017eww ; LIGOScientific:2017ikf . Due to this additional loop population, LVK analysis LIGOScientific:2021nrg resulted in stringent constraints on cosmic string tension in this scenario, limiting it to $G\mu\lesssim 4\times 10^{-15}$, and consequently LISA is only expected to allow us to probe this model in a narrow range of tensions of $G\mu\sim\mathcal{O}(10^{-15})-\mathcal{O}(10^{-16})$ Boileau:2021gbr . In this range of tensions, as was shown in Auclair:2019wcv , the effects of this additional loop population are not significant in the LISA frequency range and therefore the predictions of this model closely resemble that of the other models discussed here in the LISA window.
@@ 2.2 Template based on analytical modeling (Model I)
  • Loops are created with a length $l$ that is a fixed fraction $\alpha_{L}\leq 1$ of the characteristic length $L$ of the cosmic string network: $l=\alpha L$.
Under these assumptions, a description of the evolution of the energy density of the network is sufficient to fully characterize the loop distribution function throughout cosmic history and to compute the SGWB they generate.  [^8] Note that the results obtained under these assumptions may be used to construct the SGWB for any distribution of loop length at the moment of creation, as described in refs. Sanidas:2012ee ; Sousa:2020sxs . Here, as in refs. Sousa:2013aaa ; Sousa:2014gka , we use the Velocity-dependent One-scale Model Martins:1996jp to describe the evolution of the network.
In this model, the loop-size parameter $\alpha=\alpha_{L}L/t$, which characterizes the loop length at formation, is treated as a free parameter and a power spectrum of the form in eq. (3) is assumed. This model then has, besides cosmic string tension $G\mu$, three additional parameters: the loop-size parameter $\alpha$, the spectral index $q$ (assuming a power spectrum of the form of eq. (3)), and a fuzziness parameter $\mathcal{F}$. The fuzziness parameter can capture a variety of effects, such as situations in which loops are created with a distribution of lengths (and kinetic energies) instead of being created with exactly the same size Blanco-Pillado:2013qja ; Sousa:2020sxs .
The contribution of the $j$-th harmonic mode of emission to the SGWB is such that $j\mathrm{\Omega}_{\rm gw}^{j}(jf)=\mathrm{\Omega}_{\rm gw}^{1}(f)$ (cf. eq. (4)) and therefore determining the contribution of the fundamental mode of emission is sufficient to fully characterize the SGWB generated by cosmic string loops. In ref. Sousa:2020sxs , the authors derived accurate analytical approximations to the three contributions $\Omega^{r}_{\rm gw}$, $\Omega^{rm}_{\rm gw}$, and $\Omega^{m}_{\rm gw}$ in the fundamental mode of emission. Here, these will serve as the basis to construct a novel fully analytical template for Model I, $\Omega_{\rm gw}(f,q,G\mu,\alpha)$, including all relevant modes of emission in summation in eq. (2). Since this derivation is quite lengthy and technical, here we just provide a brief outline of the method used. We refer the reader to ref. Sousa:2020sxs and appendix A for a detailed derivation.
@@
where $i=r$ for $\mathrm{\Omega}_{\rm gw}^{r}$ and $\mathrm{\Omega}_{\rm gw}^{rm}$ and $i=m$ for $\mathrm{\Omega}_{\rm gw}^{m}$. Nevertheless, since computing this summation explicitly is computationally costly, we resort to the Euler–Maclaurin summation formula knopp1990theory to perform it, as detailed in Appendix A.
Here, we have also included the impact of the decrease of the number of relativistic DoF that occurs during the radiation era in $\mathrm{\Omega}_{\rm gw}^{r}$. This decrease, encoded in the function $\Delta_{r}(a)$ in (7), affects the expansion rate of the universe and thus may leave clear signatures on the spectrum. Although the change in $\Delta_{r}(a)$ is continuous, here we model it as a piecewise constant function based on the default implementation of micrOMEGAS5.0 Belanger:2018ccd , and follow the approach in ref. Sousa:2020sxs to derive an analytical correction to $\mathrm{\Omega}_{\rm gw}^{r}$ to account for this effect. Note however that, when there is a sudden change in the rate of expansion caused by a change in the number of relativistic DoF, the evolution of the string network temporarily deviates from the linear scaling regime and the network takes some time to relax back into this regime. This effect is non-linear and complex to model analytically and, for this reason, we develop a fit using the full numerical computation of the SGWB (obtained using the approach in ref. Sousa:2013aaa ). We also describe this fit in more detail in appendix A.
@@ 2.3 Template based on simulation-inferred modeling (Model II)
We characterize $P_{j}$ by what is known as the BOS spectrum Blanco-Pillado:2015ana ; Blanco-Pillado:2017oxo , instead of using the power-law approximation in (3999The literature often uses “Model II” to mean a cosmic string SGWB model which adopts the loop number densities of eq. (9) without specifying the $P_{j}$, i.e., without distinguishing between the power-law and BOS approaches. Here, by “Model II”, we will always mean the one which uses the BOS $P_{j}$. We will continue to specify this approach in the text to avoid any confusion with the literature. This spectrum is computed from a set of around $1000$ non-self-intersecting loops extracted from the scaling population of loops obtained from the cosmological network simulations Blanco-Pillado:2017oxo . The calculation of this spectrum takes into account backreaction by using a toy model based on a smoothing procedure. The result of this approach leads to a cusp-like spectrum at high $j$ and approximately matches with Model I when that model sets $\alpha=0.1$, $q=4/3$, and $\mathcal{F}=0.1$ (see Appendix A). The spectra are thus similar to those in the top left panel of figure 1.101010For a more comprehensive comparison of the spectra predicted in model I and model II we refer the reader to Auclair:2019wcv .
Creating an SGWB using Model II involves the use of data tables, most notably when taking the BOS power spectrum, but also in the DoF $\Delta_{r}(a)$, which ref. Blanco-Pillado:2017oxo indicates is an implementation courtesy of Masaki Yamada. As such, there is no closed-form expression for this $\Omega_{\rm gw}$ (even in the instance of analytic, single-index spectra such as the cusp or kink models). Instead, for the template subroutine, we make use of a comprehensive data table, plus interpolation, to determine the SGWB injected and recovered.
@@ 2.4 Effects of non-standard cosmologies on the SGWB
2.4.1 Modified equation of state
The first type of cosmological modification we consider is an early period, preceding the recent radiation dominated era, during which the universe is dominated by a new energy component, leading to a non-standard equation of state for the cosmos Allahverdi:2020bys . For example, an early matter domination era with, $\rho\propto a^{-3}$, may result from an appreciable abundance of a long-lived massive particle, oscillations of scalar fields or primordial black holes Moroi:1999zb ; Gouttenoire:2019rtn ; Bernal:2021yyb ; Ghoshal:2023sfa . Such an era ends with a reheating-like process as the long-lived species decay into SM particles. Another example is kination domination, where the cosmic energy is dominated by a scalar field evolving in a potential $V(\phi)\propto\phi^{N_{\phi}}$, which results in $\rho\propto a^{-n}$, with $n=6N_{\phi}/(N_{\phi}+2)$. In the limit of large $N_{\phi}$, $N_{\phi}\to\infty$, we have $n\to 6$ and the oscillation energy is dominated by the kinetic energy of the scalar. Kination may arise in certain theories related to inflation, quintessence, dark energy, and axions Salati:2002md ; Chung:2007vz ; Poulin:2018dzj ; Gouttenoire:2021jhk . While we will not consider these cases directly, more complicated histories including combinations of these mentioned above are also possible Gouttenoire:2021jhk ; Co:2021lkc ; Servant:2023tua . In order to retain the successful predictions of BBN theory Hannestad:2004px , these new phases in cosmic evolution must transit to the standard radiation domination before $T\sim 5~{}\rm MeV$.
Since Nambu-Goto simulations are mostly restricted to the radiation and matter eras (and Minkowski backgrounds as well), here we will resort to Model I to developed templates to describe these scenarios — that we will label as model Ia. Let us then assume that, before the recent radiation era, at temperatures $T>T_{\rm rd}$, the energy density of the universe is dominated by a component whose energy density scales as $\rho\propto a^{-3(1+w)}$, where $w$ is the equation-of-state parameter for this period. There are then two additional free parameters in this model, $T_{\rm rd}$ and $w$, besides the $3+1$ free parameters of Model I. In this scenario, for large enough frequencies $f>f_{\rm rd}$, the contribution of the fundamental mode of emission to the SGWB is modified as Cui:2017ufi ; Cui:2018rwi ; Sousa:2020sxs :